OpenStudy (kl0723):

Integral only with u-sub method:

OpenStudy (kl0723):

\[\int\limits_{0}^{1} \frac{ dx }{ (1+\sqrt{x})^4 }\] any hints?

OpenStudy (anonymous):

\[u=1+\sqrt{x},\]\[du=(1/2\sqrt{x})dx\]\[2\sqrt{x}du=dx\]\[2udu=dx\]

OpenStudy (anonymous):

and then do you limit changes, and then integrate

OpenStudy (kl0723):

looks like I have to apply double integral subsbtitution

OpenStudy (kl0723):

I did that yes :)

OpenStudy (kl0723):

I'm working on it right now

OpenStudy (anonymous):

x=1 u=2 x=0 u=1

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\frac{2u^2}{u^4}~du\]

OpenStudy (anonymous):

then you can do it, I guess.

OpenStudy (kl0723):

actually my u=sqrt(x)

OpenStudy (anonymous):

no need

OpenStudy (anonymous):

make it 1+sqrt{x}. the derivative is same, and w/o any addition in denominator

OpenStudy (anonymous):

u=1+sqrt{x} is much simpler, ain't it?

OpenStudy (kl0723):

well, looks like it's going well so far lol

OpenStudy (anonymous):

yes, \[\int\limits_{1}^{2}\frac{2u^2}{u^4}~du=\int\limits_{1}^{2}2u^{-2}~du\]

OpenStudy (anonymous):

from that point it is very simple, tell me the general anti derivative and then the antiderivative from 1 to 2.

OpenStudy (anonymous):

what is the general antiderivative?

OpenStudy (anonymous):

oh I made an error

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\frac{2u}{u^4}~du\]

OpenStudy (anonymous):

because du=1/2sqrt{x} dx 2sqrt{x} du= dx 2u du = dx

OpenStudy (anonymous):

sorry for my mistake. the u sub is good, but with one correction to my work

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\frac{2u}{u^4}~du\]

OpenStudy (kl0723):

I get that for dx

OpenStudy (kl0723):

yes, and I took the 2 out of the integral

OpenStudy (anonymous):

yes, so my posted integral is your new expression

OpenStudy (anonymous):

sure, whatever you like to do.

OpenStudy (kl0723):

I'll post when I get an answer

OpenStudy (anonymous):

\[2\int\limits_{1}^{2}u^{-3}~du\]

OpenStudy (anonymous):

\[2\int\limits_{}^{}u^{-3}~du=2(-2)u^{-2}+C\]

OpenStudy (anonymous):

go on, and beside the correct mistake, there are none mistakes

OpenStudy (kl0723):

yep, my substitution got tricky :/